OpenStudy (anonymous):

2/x-2*x^2-4/4 Please help! (Make sure that the final answer is simplified completely.)

OpenStudy (anonymous):

\[\frac{2}{x-2}\times \frac{x^2-4}{4}\] \[\frac{2(x+2)(x-2)}{4(x-2)}\] \[\frac{x+2}{2}\]